# Scaling Laws of Synthetic Data for Language Models

### v2

@@ Abstract
Large language models (LLMs) achieve strong performance across diverse tasks, largely driven by high-quality web data used in pre-training. However, recent studies indicate this data source is rapidly depleting. Synthetic data emerges as a promising alternative, but it remains unclear whether synthetic datasets exhibit predictable scalability comparable to raw pre-training data. In this work, we systematically investigate the scaling laws of synthetic data by introducing SynthLLM, a scalable framework that transforms pre-training corpora into diverse, high-quality synthetic datasets. Our approach achieves this by automatically extracting and recombining high-level concepts across multiple documents using a graph algorithm. Key findings from our extensive mathematical experiments on SynthLLM include: (1) SynthLLM generates synthetic data that reliably adheres to the rectified scaling law across various model sizes; (2) Performance improvements plateau near 300B tokens; and (3) Larger models approach optimal performance with fewer training tokens. For instance, an 8B model peaks at 1T tokens, while a 3B model requires 4T. Moreover, comparisons with existing synthetic data generation and augmentation methods demonstrate that SynthLLM achieves superior performance and scalability. Our findings highlight synthetic data as a scalable and reliable alternative to organic pre-training corpora, offering a viable path toward continued improvement in model performance.
Figure 1: Scaling laws on different model sizes. The x axis denotes the number of training tokens. The y axis represents the models’ error rates on MATH. The green points represent the data sizes used to fit the scaling laws, while the red points are used to test the prediction performance of the fitted curves.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of real-world applications, including chat assistant [39, 4] and code generation [9, 3]. Their success is fundamentally driven by the availability of vast amounts of high-quality data, which fuels model development at scale. As an essential resource, pre-training data provides the fundamental knowledge and generalization capability that underpin LLMs’ broad applicability [41, 7]. The importance of pre-training data is further reinforced by scaling laws [20, 18, 38], which suggest that larger models trained on more data tend to exhibit predictable and consistent performance improvements.
@@
However, recent studies [48, 37, 44] have raised concerns that our supply of high-quality pre-training data is rapidly depleting. With the web-scraped textual corpora that power LLMs approaching saturation, the conventional scaling paradigm may face diminishing returns, potentially slowing down future progress. Rather than continuing to rely solely on next-token prediction over raw pre-training data, it is crucial to explore more principled methods to better harness the potential of existing corpora. One promising direction is the use of synthetic data [13, 27, 30, 35, 32, 1], which offers a way to amplify and extend the utility of existing human-generated data, potentially sustaining performance improvements even as natural data resources dwindle. This prompts a natural question: Are there scaling laws for synthetic data? While scaling laws for pre-training data are well-documented [20, 18], whether similar principles apply to synthetic data remains unclear. Can scaling synthetic datasets sustain performance gains, or do fundamental limitations arise? Understanding its scaling behavior is key to assessing its long-term viability as a solution to data scarcity.
@@
To investigate the scaling laws of synthetic data, we aim to design a scalable approach capable of generating synthetic data at the scale. Conventional approaches to synthetic dataset curation depend heavily on limited human-annotated seed examples from target domains [50, 43, 53, 36, 46, 12, 23]. This dependency fundamentally constrains both the diversity and scalability of the resulting datasets. In contrast, pre-training corpus—being both vast and highly diverse—still remains an underutilized resource for scalable synthetic data generation. We develop SynthLLM, a scalable web-scale synthetic data generation method designed to systematically transform pre-training data into high-quality synthetic datasets.
@@
SynthLLM operates through three distinct stages. Initially, our method autonomously identifies and filters high-quality web documents within a target domain (e.g., Mathematics). Subsequently, leveraging these high-quality reference documents, we generate large-scale, diverse questions (or prompts) by employing open-source LLMs through three complementary methods, each strategically designed to progressively enhance question diversity. In final stage, we produce corresponding answers (or responses) to these generated questions, again utilizing open-source LLMs. It is worth noting that in the second stage, previous approaches have typically employed either direct question extraction approaches [55, 56, 54] or implemented document back-translation [26] to generate questions. However, these methods demonstrate inherent limitations in scalability, as question generation is constrained either by the finite size of reference documents containing high-quality questions or by the necessity of training dedicated back-translation models. Our method goes beyond direct extraction by automatically extracting and randomly combining high-level concepts from multiple documents with a graph algorithm, while establishing grounding across reference documents. Detailed ablation studies highlight the superior performance and scalability gained by our methods.
@@
We apply SynthLLM to the mathematical reasoning domain to investigate scaling laws in synthetic data. We observe the following: (1) Synthetic data consistently follow the rectified scaling law [29] across various model sizes, as illustrated in Figure 1. (2) Performance gains start to diminish once the amount of synthetic data exceeds approximately 300B tokens. (3) Larger models reach optimal performance more quickly compared to smaller ones. For instance, the 8B model requires only 1T tokens to achieve its best performance, whereas the 3B model needs 4T tokens.
These results underscore the critical importance of scaling synthetic data. Even with limited pre-training data, systematically expanding the synthetic dataset yields sustained and predictable performance gains, reinforcing its role as a scalable solution to reduce reliance on human-generated data while enhancing model capabilities.
2 Related Work
Synthetic Data Generation.
Synthetic data has emerged as a promising approach to augment real data, addressing challenges related to data scarcity and the high cost of data collection and annotation. It has been widely applied in pre-training, instruction following, and reasoning tasks, demonstrating notable improvements in model performance. In this work, we focus on synthetic data generation for the post-training phase. Existing synthetic data generation methods typically rely on a small set of high-quality human-annotated seed samples, leveraging LLMs to generate diverse augmentations through various techniques: 1) sampling seed instructions as few-shot examples to elicit LLMs to generate new instructions [50, 19, 47, 23]; 2) leveraging LLMs to rephrase seed samples [53, 35, 12, 21, 34]. However, the scarcity of high-quality seed examples limits the scalability and diversity of generated data. To address this, Xu et al. [51] and Li et al. [24] propose generating synthetic data from scratch, leveraging either the inherent uncertainty of LLMs or a predefined knowledge taxonomy to guide the generation process. In contrast, pre-training data—being both vast and highly diverse—still remains an underutilized resource for scalable post-training data generation. Recent methods have explored extracting high-quality samples directly from web documents (similar to our Level-1 method) [55, 56, 54] or employing document backtranslation to generate questions [26]. Our SynthLLM aligns with this line of research and offers a more efficient approach to leveraging web documents for scaling diverse sample generation compared to aforementioned methods.
Scaling Law of LLMs.
Scaling laws provide a predictive framework for estimating model performance based on key factors such as model size and pre-training data size, and have been extensively studied [17, 42, 20, 18, 5]. These laws offer valuable insights into how performance scales with computational resources, enabling more informed decisions on the optimal allocation of compute for pre-training LLMs [20, 18]. Recently, more fine-grained scaling laws have been proposed, including data-constrained scaling [37], hyperparameter scaling [6], and model distillation [8]. For LLM fine-tuning, Hernandez et al. [16] investigate scaling laws in the transition from unsupervised pre-training to fine-tuning, highlighting the critical role of pre-training in altering the scaling behavior of the model. Lin et al. [29] propose a rectified scaling law tailored specifically for fine-tuning LLMs on downstream tasks. Our work extends the rectified scaling law [29] and is the first to investigate and empirically validate that scaling laws hold for fine-tuning language models with synthetic data.
@@ 3 Scaling Law of Synthetic Data
Kaplan et al. [20] demonstrates that the performance of LLMs exhibits a power-law relationship with respect to both model size and dataset size, primarily focusing on the pre-training phase utilizing organic data. In this section, we investigate the scaling properties of synthetic data. Specifically, we aim to determine whether a scaling law exists for synthetic data that enables accurate prediction of the performance of LLMs trained on such data. Details of the approach employed for synthetic data generation (i.e., SynthLLM) will be presented in Section 4, which is a method for curating synthetic data from carefully filtered documents drawn from the pre-training corpus (comprised exclusively of organic raw data).
@@ 3.1 Scaling Laws for Organic Data
The scaling law for pre-training on organic data has been extensively studied in the literature [20, 18]. In particular, the predictive loss is a parametric function of the model parameter size $N$ and data size $D$ (i.e., number of tokens) following the power-law form [18]:
@@
This function above models the relationship between scaling loss $L(N,D)$ (usually measured on a validation set), the number of model parameters $N$, and the number of training tokens $D$, based on the classical risk decomposition, where the final loss can be decomposed into multiple terms, each corresponding to a different source of error. $L(N,D)$ is the final scaling loss. $N$ represents parameter count of the LLM, and $D$ denotes the number of training tokens. $E$, the irreducible term, reflects the fundamental loss inherent in an ideal generative process operating over the natural text distribution. $\alpha$ and $\beta$ are decay exponents that determine how the loss decreases as $N$ and $D$ increase. $A$ and $B$ regulate the rate of loss convergence and are usually influenced by the model architecture and the characteristics of the data distribution. The coefficients $\alpha,\beta,A,B,$ and $E$ are learnable parameters that are determined by fitting a scaling curve.
Hernandez et al. [16] investigate scaling laws for transfer (i.e., the transition from unsupervised pre-training to fine-tuning). Their findings indicate that fine-tuning on pre-trained models exhibits more data efficient scaling behavior compared to training models from scratch on downstream tasks, emphasizing the crucial influence of pre-training on scaling dynamics. Extending this line of inquiry, Lin et al. [29] analyze scaling laws during fine-tuning of LLMs and propose Rectified Scaling Law:
@@
In the following, we illustrate how the rectified scaling law is obtained from the vanilla scaling law in Eq. (1).
Compared to the original formulation in Eq. (1), we first remove the term associated with model parameters, as the LLM remains fixed during fine-tuning. Consequently, we derive the following marginalized form of Eq. (1):
@@
Here, $L(D)$ represents the validation metric used to evaluate the LLM’s performance on the downstream task. The irreducible term $E$ empirically represents the optimal achievable model performance as the data size approaches infinity. This term depends not only on inherent loss of the data distribution but also on the capacity limitations of the model itself.
Compared to the formulation in Eq. (3), the Rectified Scaling Law introduces an additional term, $D_{l}$, which represents the pre-learned data size. This term quantifies the latent knowledge relevant to the downstream task that the model has already acquired during pre-training. Fine-tuning on a pretrained LLM often yields better performance than training from scratch on a downstream task. The inclusion of $D_{l}$ reflects this benefit. Consequently, we obtain the final formulation in Eq. (2). The term $\frac{B}{D_{l}}$ essentially defines the initial performance of the pre-trained LLM. Lin et al. [29] show that this modification offers a more precise representation of the impact of pre-training on the model performance in downstream tasks during fine-tuning.
@@ 3.2 Scaling Laws for Synthetic Data
Prior research [20, 18, 16, 29] has primarily focused on investigating the scaling properties of organic, real-world data and their impact on LLM performance. To our knowledge, this work is the first to systematically investigate and empirically validate that analogous scaling laws hold for synthetic data. We employ SynthLLM (Section 4) to generate data from a carefully filtered subset of the pre-training corpus, ensuring both scale and diversity (see Section 5.3 for details). Using the full synthetic dataset from the three methods in SynthLLM, we conduct continued training on LLMs of varying sizes (Llama-3.2-1B, Llama-3.2-3B, and Llama-3.1-8B) with progressively larger subsets (100K, 200K, 400K, 800K, 1.6M, 3.2M, and 7.4M examples). We selected the mathematical reasoning domain for our experiments due to its well-established evaluation protocols and metrics. Specifically, we report the error rate on the MATH dataset [15], which covers a wide range of mathematical subjects and incorporates varied difficulty levels (ranging from Level-1 to level 5), making it more comprehensive than alternative datasets. Note that we report error rate rather than accuracy to maintain consistency with previous scaling law literature [20, 18]. It is worth noting that both lower loss and lower error rate indicate superior performance in evaluations. See more details of our synthetic datasets and experiments in Section 5.1.
Overall Trend.
@@
As shown in Figure 1, the synthetic data generated by SynthLLM adheres to the rectified scaling law (Equation 2). The $x$-axis represents the synthetic dataset size, measured in number of tokens, while the $y$-axis depicts the error rate on the MATH dataset. Both axes employ logarithmic scaling, consistent with previous work [20]. The green points denote the data sizes utilized to fit the scaling laws, whereas the red points serve to evaluate the predictive accuracy of the fitted curves. Our results demonstrate that synthetic data generated by SynthLLM consistently adheres to the rectified scaling law across various model sizes.
In addition to the rectified scaling law (Equation 2), we also attempted to fit our results with the power scaling law (Equation 3). Our analysis reveals that the power law fails to capture the scaling trend effectively (see Section 5.4). We hypothesize that this phenomenon may be attributed to the base models used in our experiments having already acquired certain mathematical problem-solving capabilities during their pre-training phase.
Model Size Matters.
The decay exponent $\beta$ in Eq. (2) quantifies the rate at which the error rate decreases as dataset size increases. A larger value of $\beta$ corresponds to a more rapid reduction in error rate, given the same amount of synthetic tokens. As in Figure 1, in comparison to its large counterparts (3B and 8B models), the 1B model exhibits a smaller $\beta$ value ($0.34$), suggesting that increasing the volume of synthetic data yields more modest performance improvements for this smaller model. The ratio $\frac{B}{D_{l}}$ demonstrates a strong correlation with the initial capacity of models, wherein a smaller value (smaller error rate) indicates greater initial model capability. The 7B model exhibits the smallest ratio value ($56.3$),
@@
which may be attributed to its superior model capacity and more comprehensive pre-training data exposure. Furthermore, the 3B and 7B models demonstrate identical decay exponents ($\beta$), indicating that performance disparities stem predominantly from differences in initial knowledge acquisition and inherent model capacity rather than from variations in the efficacy of synthetic data scaling.
Figure 2: Scaling laws on different model sizes. The x-axis represents the number of training tokens. The y-axis shows the models’ error rates on MATH benchmark. Solid lines indicate fitted scaling curves for different model sizes. The dotted lines represent predicted final performances when synthetic data is further scaled, as extrapolated from these curves. More details are shown in Figure 1.
Predictable Performance.
For 3B and 8B models, we evaluated the predictive capability of our fitted scaling curves. As illustrated by the red squares in Figure 1, these curves accurately forecast error rates on the MATH benchmark for both models ($40.0\%$ and $28.7\%$, respectively), when synthetic dataset is expanded to 7.4M samples (4.5B tokens). These error rates correspond to accuracies of $60.0\%$ and $71.3\%$. This highlights the critical role of scaling synthetic data, demonstrating that even as pre-training data nears saturation, systematically expanding the synthetic dataset continues to yield sustained and predictable performance gains. Leveraging our established scaling curves, we can extrapolate the future performance of models with continued expansion of synthetic data beyond current volumes. Table 1 presents these projections for the 3B and 8B models. The results indicate that performance improvements begin to diminish after approximately 300B synthetic tokens. We estimate that the 3B model would require approximately 4T synthetic data tokens to approach its performance ceiling, while the 8B model would need approximately 1T synthetic tokens (also see Figure 2).

### v3

@@ Abstract
Large language models (LLMs) achieve strong performance across diverse tasks, driven by high-quality web data used in pre-training. However, recent studies indicate web data is rapidly depleting. Synthetic data emerges as a promising alternative, but it remains unclear whether synthetic datasets exhibit predictable scalability comparable to raw pre-training data. In this work, we systematically investigate scaling laws of synthetic data by introducing SynthLLM, a scalable framework that transforms pre-training corpora into diverse, high-quality synthetic datasets. Our approach achieves this by automatically extracting and recombining high-level concepts across multiple documents using a graph algorithm. Key findings from our experiments with SynthLLM on math domain include: (1) SynthLLM generates synthetic data that reliably adheres to rectified scaling law across various model sizes; (2) Performance gains gradually diminish near 300B tokens; and (3) Larger models approach optimal performance with fewer training tokens. For instance, an 8B model peaks at 1T tokens, while a 3B model requires 4T. Moreover, comparisons with existing synthetic data generation and augmentation methods demonstrate that SynthLLM achieves superior performance and scalability. Our findings highlight synthetic data as a scalable and reliable alternative to raw pre-training data, offering a viable path toward continued improvement in model performance.
@@ 1 Introduction
Large language models (LLMs) achieve remarkable capabilities, primarily driven by vast amounts of high-quality pre-training data, which provide the foundational knowledge and generalization capacity necessary for their broad applicability (Radford et al., 2019; Brown et al., 2020). This relationship is further reinforced by scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022; Muennighoff et al., 2023), indicating that increasing both model size and data volume typically yields predictable and consistent performance improvements.
@@
However, recent studies (Villalobos et al., 2024; Muennighoff et al., 2023) indicate that the supply of high-quality pre-training data is rapidly diminishing. As web-scraped textual corpora approach saturation, conventional scaling methods may face diminishing returns, potentially slowing future progress. Therefore, it becomes crucial to explore more principled approaches that better harness the potential of existing corpora. Synthetic data (Gunasekar et al., 2023; Li et al., 2023; Liu et al., 2024; Maini et al., 2024; Long et al., 2024; Abdin et al., 2024) emerges as a promising alternative, providing a means to amplify and extend the utility of human-generated data. Yet, it remains unclear whether synthetic data follows predictable scaling laws akin to those established for natural pre-training data (Kaplan et al., 2020; Hoffmann et al., 2022). This raises a critical question: Can synthetic data sustain scalable performance improvements, or are there fundamental limitations? Understanding this is essential to assessing the viability of synthetic data as a long-term solution to data scarcity.
@@
To investigate the scaling laws of synthetic data, we aim to develop a scalable approach for generating synthetic data at the scale. Conventional synthetic dataset generation methods rely heavily on limited human-annotated seed examples from target domains (Wang et al., 2022; Tang et al., ; Yu et al., 2024; Mitra et al., 2023; Toshniwal et al., 2025; Ge et al., 2024; Li et al., 2024a), fundamentally restricting the diversity and scalability of the resulting datasets. In contrast, pre-training corpora—vast and inherently diverse—remain an underutilized resource for synthetic data generation. To address this gap, we introduce SynthLLM, a scalable framework designed to systematically transform pre-training data into high-quality synthetic datasets.
@@
SynthLLM operates through three distinct stages. First, it autonomously identifies and filters high-quality web documents from the target domain (e.g., mathematics). Second, using these selected reference documents, it generates diverse, large-scale questions or prompts through open-source LLMs employing three complementary methods designed to progressively enhance diversity. Finally, the method produces corresponding answers to these questions, again utilizing open-source LLMs. Previous approaches typically used direct question extraction (Yue et al., 2024; Zhou et al., 2024; Yuan et al., 2025) or document back-translation (Li et al., 2024). However, these methods have inherent scalability limitations, constrained by either the finite number of reference documents containing quality questions or the necessity of dedicated back-translation model training. SynthLLM goes beyond direct extraction by automatically extracting and randomly combining high-level concepts from multiple documents with a graph algorithm, while establishing grounding across diverse reference documents. Detailed ablation studies highlight the superior performance and scalability gained by our methods.
@@
We apply SynthLLM to math reasoning domain to study the scaling laws of synthetic data. We observe the following: (1) Synthetic data consistently follow rectified scaling law (Lin et al., 2024) across various model sizes, as illustrated in Figure 1. (2) Performance gains start to diminish once the amount of synthetic data exceeds approximately 300B tokens. (3) Larger models reach optimal performance more quickly compared to smaller ones. For instance, the 8B model requires only 1T tokens to achieve its best performance, whereas the 3B model needs 4T tokens. These results underscore the critical importance of scaling synthetic data. Even with limited pre-training data, systematically expanding the synthetic dataset yields sustained and predictable performance gains, reinforcing its role as a scalable solution to reduce reliance on human-generated data while enhancing model capabilities.
@@ 2 Scaling Law of Synthetic Data
In this section, we first investigate and demonstrate the scaling properties of synthetic data. Details of the approach employed for synthetic data generation (i.e., SynthLLM) will be presented in Section 3.
@@ 2.1 Scaling Laws for Organic Data
The scaling law for pre-training on organic data has been extensively studied in the literature (Kaplan et al., 2020; Hoffmann et al., 2022). In particular, the predictive loss is a parametric function of the model parameter size $N$ and data size $D$ (i.e., number of tokens) following the power-law form (Hoffmann et al., 2022):
@@
The function above models the scaling loss L(N,D)L(N,D)—typically measured on a validation set—as a relationship between the number of model parameters $N$ and training tokens $D$, based on classical risk decomposition. This decomposition splits the final loss into multiple terms, each corresponding to a distinct source of error. Here, $N$ denotes the LLM parameter count, and $D$ represents the number of training tokens. The irreducible term $E$ reflects the minimal achievable loss inherent to the ideal generative process over natural text distributions. The decay exponents $\alpha$ and $\beta$ control how rapidly the loss diminishes as $N$ and $D$ increase, respectively. The coefficients $A$ and $B$, typically influenced by model architecture and data characteristics, regulate the rate of loss convergence. All parameters $\alpha,\beta,A,B,$ and $E$ are learned by fitting the scaling curve to empirical data.
Hernandez et al. (2021) explore scaling laws for transfer learning, specifically the transition from unsupervised pre-training to fine-tuning. Their results indicate that fine-tuning pretrained models is more data-efficient than training from scratch, highlighting the importance of pre-training in shaping scaling behavior. Building upon this, Lin et al. (2024) further analyze scaling laws in fine-tuning scenarios and introduce the Rectified Scaling Law:
@@
In the following, we illustrate how the rectified scaling law is obtained from the vanilla scaling law in Eq. (1). Compared to the original formulation in Eq. (1), we first remove the term associated with model parameters, as the LLM remains fixed during fine-tuning. Consequently, we derive the following marginalized form of Eq. (1):
@@
Here, $L(D)$ denotes the validation metric for evaluating the LLM’s performance on downstream tasks. The irreducible term $E$ represents the optimal performance achievable by the model as data size approaches infinity, reflecting both inherent data-distribution loss and model capacity constraints. Compared to Eq. (3), the Rectified Scaling Law introduces an additional parameter, $D_{l}$, termed the pre-learned data size. This quantifies the latent knowledge relevant to the downstream task acquired during pre-training. Including $D_{l}$ captures the empirical observation that fine-tuning pretrained models yields better performance than training from scratch. Consequently, we obtain the final formulation in Eq. (2). Specifically, the term $\frac{B}{D_{l}}$ characterizes the initial performance advantage provided by pre-training. Lin et al. (2024) demonstrate that this refinement provides a more accurate depiction of how pre-training influences downstream task performance during fine-tuning.
@@ 2.2 Scaling Laws for Synthetic Data
Prior studies (Kaplan et al., 2020; Hoffmann et al., 2022; Hernandez et al., 2021; Lin et al., 2024) have primarily investigated scaling laws of organic data. To our knowledge, this is the first systematic exploration and empirical validation of scaling laws for synthetic data. We employ SynthLLM (Section 3) to generate data from a carefully filtered subset of the pre-training corpus, ensuring both scale and diversity (see Section 4.3 for details). Using the full synthetic dataset from the three methods in SynthLLM, we conduct continued training on LLMs of varying sizes (Llama-3.2-1B, Llama-3.2-3B, and Llama-3.1-8B) with progressively larger subsets (100K, 200K, 400K, 800K, 1.6M, 3.2M, and 7.4M examples). In this work, we main focus on the mathematical reasoning domain due to its well-established evaluation protocols and metrics. Additionally, we conduct preliminary experiments in the coding domain, and present detailed results in Section 4.4. Model performance is assessed by reporting error rates on the MATH dataset (Hendrycks et al., 2021), which covers a wide range of mathematical subjects and difficulty levels (1–5). We use error rate as the evaluation metric, where lower error rates indicate better performance. Further details on datasets and experiments are provided in Section 4.1.
Figure 1: Scaling laws across different model sizes. The x-axis denotes the number of training tokens, while the y-axis shows error rates on the MATH benchmark. Green points indicate data sizes used for fitting the scaling curves, and red points assess the predictive accuracy of these fitted curves.
@@
Overall Trend. As shown in Figure 1, synthetic data generated by SynthLLM adheres to the rectified scaling law (Equation 2). The $x$-axis represents the synthetic dataset size, measured in number of tokens, while the $y$-axis depicts the error rate on the MATH dataset. Both axes employ logarithmic scaling, consistent with previous work (Kaplan et al., 2020). The green points denote the data sizes utilized to fit the scaling laws, whereas the red points serve to evaluate the predictive accuracy of the fitted curves. We additionally attempted fitting our data using Equation 3, but found that it fails to accurately model our results (see Section C.2). We hypothesize this discrepancy arises because the base models employed in our experiments already possess certain mathematical problem-solving capabilities acquired during pre-training.
Model Size Matters The decay exponent $\beta$ in Eq. (2) quantifies how quickly the error rate decreases as dataset size grows; a larger $\beta$ indicates a more rapid improvement given same synthetic token volume. As shown in Figure 1, 1B model exhibits a smaller $\beta$ ($0.34$) compared to larger models (3B and 8B), suggesting it benefits less from increased synthetic data. The ratio $\frac{B}{D_{l}}$ correlates strongly with initial model performance, where a smaller ratio implies higher initial capability. The 7B model achieves the lowest ratio ($56.3$), reflecting its superior model capacity and richer pre-training exposure. Besides, identical decay exponents ($\beta$) for 3B and 7B models indicate that performance differences primarily arise from disparities in initial knowledge and model capacity rather than synthetic data scaling efficacy.
Predictable Performance We evaluated the predictive accuracy of our fitted scaling curves for the 3B and 8B models. As shown by the red squares in Figure 1, these curves reliably forecast the models’ error rates (40.0% for 3B, 28.7% for 8B) on the MATH benchmark when scaled to 4.5B synthetic tokens, corresponding to accuracies of 60.0% and 71.3%, respectively. These results underscore the effectiveness of scaling synthetic data: even as traditional pre-training data approaches saturation, systematic expansion of synthetic datasets continues to provide stable and predictable performance improvements. Using our scaling curves, we further extrapolate the future performance of the 3B and 8B models as the volume of synthetic data continues to grow. Table 1 presents these projections, indicating diminishing returns in performance improvements beyond approximately 300B synthetic tokens.
@@
Specifically, we estimate the 3B model would require around 4T synthetic tokens to approach its performance ceiling, whereas 8B model would need about 1T tokens. Currently, SynthLLM generates only five questions per document from pre-training corpus, which is still far from fully leveraging the potential of each document (see Figure 5). We can easily continue scaling up our synthetic dataset to further approach the performance ceiling.